OpenStudy (anonymous):

A ball is dropped from the top of a 54m high cliff. At the same time, a carefully aimed stone is thrown straight up from the bottom of the cliff with a speed of 27m/s^2 . The stone and ball collide part way up.

OpenStudy (anonymous):

How far above the base of the cliff does this happen?

OpenStudy (anonymous):

If I were to make an equation for the height of the dropped ball as a function of time, it would be \[ h(t) = H -\frac{1}{2} gt^2 \] where H is the height of the cliff. Can you make an equation for the height of the ball that was thrown upward?

OpenStudy (anonymous):

i have \[y=\frac{ v ^{2}-v _{0}^{2} }{ 2g }\] \[v ^{2}=27m/s^2\]

OpenStudy (anonymous):

wait, i got that wrong. hold on

OpenStudy (anonymous):

As a function of time, the height from the bottom of the cliff can be written \[ h(t) = v_0 t - \frac{1}{2}gt^2 \]

OpenStudy (anonymous):

Set those two equal to one another. What do you get (solving for t) ?

OpenStudy (anonymous):

i got 2 seconds

OpenStudy (anonymous):

Symbolically, \[t = \frac{H}{v_0} \] Plug that into either of the two height formulas, and what do you get?

OpenStudy (anonymous):

so they collide at 34.4 meters

OpenStudy (anonymous):

\[h(t) = H - \frac{1}{2}gt^2 = H - \frac{gH^2}{2v_0^2} \] or \[h(t) = v_0t - \frac{1}{2}gt^2 = H - \frac{gH^2}{2v_0^2} \]

OpenStudy (anonymous):

sweet. thank you for your help again, you are the man. i have to go to class now but ill see you around here again. take care

OpenStudy (anonymous):

No problem. Likewise.